OpenStudy (mathisfun13):

The graph of this equation will be which conic section? a.hyperbola b.circle c.ellipse d.parabola

OpenStudy (anonymous):

all

OpenStudy (mathisfun13):

it cant be...

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Conic_section take a look at a photo on this page

OpenStudy (mathisfun13):

thats okay

ganeshie8 (ganeshie8):

wheres the equation ?

OpenStudy (mathisfun13):

lol oh . y2-x=20

ganeshie8 (ganeshie8):

\(y^2 -x = 20\)

ganeshie8 (ganeshie8):

like this ?

OpenStudy (mathisfun13):

yes

ganeshie8 (ganeshie8):

\(x = y^2-10\)

ganeshie8 (ganeshie8):

its a quadratic in y, so it gives a parabola.

ganeshie8 (ganeshie8):

you may graph it and see, plugin \(x = y^2-10\) in wolfram